3.Find the Total revenue, average revenue and marginal revenue as the total revenue equation is given as;

TR= 100-10Q^2 , from Q=0 to Q=6.


1
Expert's answer
2021-02-23T14:33:20-0500

(a) Let's calculate the TR function for each quantity:


"TR_0=100-10\cdot0^2=\$100,""TR_1=100-10\cdot1^2=\$90,""TR_2=100-10\cdot2^2=\$60,""TR_3=100-10\cdot3^2=\$10,""TR_4=100-10\cdot4^2=-\$60,""TR_5=100-10\cdot5^2=-\$150,""TR_6=100-10\cdot6^2=-\$260."


(b) Then, we can calculate the average revenue for each quantity:


"AR_1=\dfrac{TR_1}{Q_1}=\dfrac{\$90}{2}=\$45,""AR_2=\dfrac{TR_2}{Q_2}=\dfrac{\$60}{3}=\$20,""AR_3=\dfrac{TR_3}{Q_3}=\dfrac{\$10}{3}=\$3.3,""AR_4=\dfrac{TR_4}{Q_4}=\dfrac{-\$60}{4}=-\$15,""AR_5=\dfrac{TR_5}{Q_5}=\dfrac{-\$150}{5}=-\$30,""AR_6=\dfrac{TR_6}{Q_6}=\dfrac{-\$260}{6}=-\$45.3."

c) In order to find MR we need to take derivative of TR with respect to Q:

"MR=\dfrac{dTR}{dQ}=-20Q."


Let's calculate MR for each quantity:


"MR_1=\dfrac{TR_1-TR_0}{Q_1-Q_0}=\dfrac{\$90-\$100}{1-0}=-\$10,""MR_2=\dfrac{TR_2-TR_1}{Q_2-Q_1}=\dfrac{\$60-\$90}{2-1}=-\$30,""MR_3=\dfrac{TR_3-TR_2}{Q_3-Q_2}=\dfrac{\$10-\$60}{3-2}=-\$50,""MR_4=\dfrac{TR_4-TR_3}{Q_4-Q_3}=\dfrac{-\$60-\$10}{4-3}=-\$70,""MR_5=\dfrac{TR_5-TR_4}{Q_5-Q_4}=\dfrac{-\$150-(-\$60)}{5-4}=-\$90,""MR_6=\dfrac{TR_6-TR_5}{Q_6-Q_5}=\dfrac{-\$260-(-\$150)}{6-5}=-\$110."